OpenStudy (anonymous):

Rationalise the denominator A) 7/root2+root2 B)5+root2/5-root2 C)root3+root5/root3-root5

OpenStudy (australopithecus):

multiply the top by the conjugate. so for \[\frac{7}{\sqrt{2} + \sqrt{2}}\] multiple the top and the bottom by \[\sqrt{2} - \sqrt{2}\]

OpenStudy (australopithecus):

when I say the top and bottom I mean the numerator and denomoniator

OpenStudy (australopithecus):

try it, if you have any problems just post them and I will try to assist you

OpenStudy (australopithecus):

Also as an aside always use the conjugate from the denominator